OpenStudy (anonymous):

Find the mean deviation about mean for the series a , a+d , a+2d , a+3d ......... , a+ nd ?

terenzreignz (terenzreignz):

Mean is just average, right? So it's just \[\huge \frac{\sum_{k=0}^{n}a+kd}{n+1}\] Unless I'm missing something?

OpenStudy (anonymous):

\[A=\frac{ a+b }{ 2}\]

terenzreignz (terenzreignz):

Now... \[\huge \sum_{k=0}^{n}a+kd = \sum_{k=1}^{n+1}a+(k-1)d=\sum_{k=1}^{n+1}a-d+kd\]

terenzreignz (terenzreignz):

\[\huge =\sum_{k=1}^{n+1}(a-d) + \sum_{k=1}^{n+1}kd\]

terenzreignz (terenzreignz):

\[\huge \sum_{k=1}^{n+1}(a-d)+ d\sum_{k=1}^{n+1}k\]

terenzreignz (terenzreignz):

Using a rather known formula for the second summation, we get \[\huge (a-d)(n+1) + \left(\frac{d}{2}\right)(n+1)(n+2)\]

terenzreignz (terenzreignz):

dividing by n+1 to get the mean, we get \[\huge a - d + \frac{d}{2}(n+2)\] And that's the average, I reckon :)